OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

OpenStudy (irishboy123):

\[v • w = <v_x, v_y, v_z>•<w_x,w_y, w_z> = v_x w_x + v_y w_y + v_z w_z\]

OpenStudy (anonymous):

r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6> okay so <-3,-8,-3>* <-4,-2,-6>= 3x-4x+-8y8y+-3z-6 ?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

let's do r • r r = <8, 8, -6> r • r = <8, 8, -6>•<8, 8, -6> = <8*8 + 8*8 + (-6)*(-6)> = 64 + 64 + 36 = 164 now, let's do r•v r•v = <8, 8, -6>•<3, -8, -3> = 8*3 + 8*(-8) + (-6)*(-3) = 24 -64 + 18 = -22 do you get the idea? r, v & w are vectors. the "dot" product turns vectors into simple (scalar) numbers. why? what is the point of it all? well, the idea is incredibly powerful as the dot product is a black box that contains the Law of Cosines without the need for all of the usual kerfuffle -- and it make life incredibly easy in 3-D. however, you have to be able to process it first. just churn the numbers.

OpenStudy (anonymous):

ohh okayyy it makes sense now so -22 would be the answer!?

OpenStudy (anonymous):

thank you so much!!!! @IrishBoy123

OpenStudy (irishboy123):

v = <3, -8, -3>; w = <-4, -2, -6> v•w = <3, -8, -3>•<-4, -2, -6> = 3(-4) + (-8)(-2) + (-3)(-6) = -12 +16 + 18 = +22 have i confused you, or got the sums wrong. ie, where does the minus sign come into it?

OpenStudy (anonymous):

ohh i see so what would i do now! just calculate the last part? -12 +16 + 18 = +22? @IrishBoy123

OpenStudy (anonymous):

oh wait nevermind lol the answers 22!

OpenStudy (irishboy123):

+1

OpenStudy (anonymous):

wait what lol

OpenStudy (irishboy123):

+1 means "well done" that's all.

OpenStudy (anonymous):

oh ahahahaha thanks!! youre a lifesaver @IrishBoy123